OpenStudy (anonymous):

What are the x-intercepts of y = 2x2 - 3x - 20?

OpenStudy (anonymous):

\[y=2x^2-3x-20\] \[when\ y=0:\] \[2x^2-3x-20=0\] \[(2x+5)(x-4)=0\] \[2x+5=0 \rightarrow x=-5/2\] \[x-4=0 \rightarrow x=4\] therefore, x-ints are (-5/2,0) and (4,0)